# Implicit Bias of Mirror Descent for Shallow Neural Networks in Univariate Regression

### v1

@@ Abstract
We examine the implicit bias of mirror flow in univariate least squares error regression with wide and shallow neural networks. For a broad class of potential functions, we show that mirror flow exhibits lazy training and has the same implicit bias as ordinary gradient flow when the network width tends to infinity. For ReLU networks, we characterize this bias through a variational problem in function space. Our analysis includes prior results for ordinary gradient flow as a special case and lifts limitations which required either an intractable adjustment of the training data or networks with skip connections. We further introduce scaled potentials and show that for these, mirror flow still exhibits lazy training but is not in the kernel regime. For networks with absolute value activations, we show that mirror flow with scaled potentials induces a rich class of biases, which generally cannot be captured by an RKHS norm. A takeaway is that whereas the parameter initialization determines how strongly the curvature of the learned function is penalized at different locations of the input space, the scaled potential determines how the different magnitudes of the curvature are penalized.
@@ 1 Introduction
Mirror descent is a broad class of first-order optimization algorithms that generalize ordinary gradient descent (Nemirovskij & Yudin, 1983). It can be viewed as using a general distance-like function, defined by the choice of a convex potential function, instead of the usual Euclidean squared distance to determine the update direction in the search space (Beck & Teboulle, 2003). The choice of the geometry of the parameter space of a learning system is crucial, as it can affect the speed of convergence and the implicit bias of parameter optimization (Neyshabur et al., 2017). The implicit bias of mirror descent has been studied theoretically for linear models (Gunasekar et al., 2018; Sun et al., 2023; Pesme et al., 2024), matrix sensing (Wu & Rebeschini, 2021), and also for nonlinear models at least in a local sense (Azizan et al., 2022). However, we are not aware of any works characterizing the implicit bias of mirror decent for neural networks describing more than the possible properties of the parameters also the nature of the returned solution functions. This is the gap that we target in this article. We focus on the particular setting of mirror flow applied to univariate least squares error regression with shallow and infinitely wide neural networks.
@@
We show for networks with standard parametrization that, under a broad class of potential functions, mirror flow displays lazy training, meaning that even though the loss converges to zero the parameters remain nearly unchanged. We show in turn that mirror flow has the same implicit bias as gradient flow when the network width tends to infinity. This is surprising because, in contrast, for finite-dimensional parameter models mirror flow has been observed to exhibit a different implicit bias than gradient descent (see, e.g., Gunasekar et al., 2018; Sun et al., 2023). We then proceed to characterize the particular solution function that is returned by mirror flow for wide ReLU networks, by expressing the difference between the network functions before and after training, $g(x)=f_{\mathrm{trained}}(x)-f_{\mathrm{initial}}(x)$, as the solution to a variational problem. By examining the function representation cost, we show that mirror flow is biased towards minimizing a weighted $L_{2}$ norm of the second-order derivative of the difference function, $\int p^{-1}(x)(g^{\prime\prime}(x))^{2}\mathrm{d}x$, where the weight $p^{-1}$ is determined by the network parameter initialization, and two additional terms that control the first-order derivative. This includes as a special case a previous result for gradient flow and ReLU activations by Jin & Montúfar (2023). Our analysis also removes a key limitation of that result, which required either an intractable linear adjustment of the training data or otherwise adding skip connections to the network. We observe that other previous works on the implicit bias of gradient descent suggested that shallow ReLU networks in the kernel regime could be well-approximated by a cubic spline interpolator of the training data (Williams et al., 2019; Heiss et al., 2019; Sahs et al., 2022) but overlooked the need of linearly adjusting the training data or adding skip connections. We illustrate that even a slight modification of the training data by applying a constant shift to all training labels can lead to significantly different solutions upon training which are biased not only in terms of curvature but also slopes.
@@ 1.1 Contributions
The goal of this article is to provide insights into the implicit bias of mirror flow for univariate regression with shallow neural networks. Our contributions can be summarized as follows.
  • We show for a broad class of potential functions that, mirror flow, when applied to networks with standard parametrization, displays lazy training. We show that in this case the implicit bias of mirror flow is equivalent to that of gradient flow when the number of hidden units tends to infinity.
  • We characterize the implicit bias of mirror flow for wide and shallow ReLU networks in function space. Concretely, we express the difference of the trained and initialized network function as the solution to a variational minimization problem in function space (Theorem 2).
  • We introduce a class of scaled potentials for which we show that mirror flow, applied to networks with standard parametrization, falls in the lazy regime but not in the kernel regime. We characterize the corresponding implicit bias for infinitely wide networks and observe it is strongly dependent on the potential and thus different from ordinary gradient flow.
  • Finally, we characterize the implicit bias of mirror flow with scaled potentials in function space for wide and shallow networks with absolute value activation function (Theorem 8). Our results show that the choice of the potential biases the magnitude of the curvature of the solution function.
@@ 1.2 Related works
The implicit bias of mirror descent has been studied for under-determined linear models. In classification problems with linearly separable data, Sun et al. (2022) show that mirror descent with homogeneous potentials converges in direction to a max-margin classifier with respect to a norm defined by the potential. Pesme et al. (2024) extend this work to non-homogeneous potentials. For regression, Gunasekar et al. (2018); Azizan & Hassibi (2019) show that mirror descent is biased towards the minimizer of the loss that is closest to the initialization with respect to the Bregman divergence defined by the potential. Azizan et al. (2022) extend this result to nonlinear models, albeit placing a strong assumption that the initial parameter is sufficiently close to a solution. For matrix sensing, Wu & Rebeschini (2021) show mirror descent with spectral hypentropy potential tends to recover low-rank matrices. Li et al. (2022) connect gradient flow with commuting reparameterization to mirror flow, and show that both have identical trajectories and implicit biases. This connection is further studied by Chou et al. (2023) specifically for linear models.
@@
For a parameterized model and a given cost function, e.g., a norm, on the parameter space, the representation cost of a target function refers to the minimal cost required to implement that function. A number of works have studied the representation cost for overparameterized networks, including wide shallow ReLU networks (Savarese et al., 2019; Ongie et al., 2020), shallow ReLU networks with added linear layers (Parkinson et al., 2024), deep networks (Jacot et al., 2022), and infinitely deep networks (Jacot, 2023). While these works relate representation cost to the implicit bias of regularized training, or more precisely, to regularized empirical risk minimization (RERM), our study demonstrates that representation cost is also crucial in understanding the implicit bias of unregularized training. Moreover, instead of studying RERM in a static and optimization-free manner, we characterize the implicit bias with concrete convergence guarantees.
@@ 2 Problem Setup
We focus on fully-connected two-layer neural networks with one input unit, $n$ hidden units and one output unit, defined by:
$f(x,\theta)=\sum_{k=1}^{n}a_{k}\sigma(w_{k}x-b_{k})+d,$ (1)
where $\sigma(z)$ is a nonlinear activation function. We write $\theta=\operatorname{vec}(\mathbf{w},\mathbf{b},\boldsymbol{a},d)\in\mathbb{R}^{3n+1}$ for the vector of all parameters, comprising the input weights $\mathbf{w}=(w_{1},\ldots,w_{n})^{T}\in\mathbb{R}^{n}$, input biases $\mathbf{b}=(b_{1},\ldots,b_{n})^{T}\in\mathbb{R}^{n}$, output weights $\boldsymbol{a}=(a_{1},\ldots,a_{n})^{T}\in\mathbb{R}^{n}$, and output bias $d\in\mathbb{R}$. We assume the parameters of the network are initialized by independent samples of mutually independent random variables $\mathcal{W},\mathcal{B},\mathcal{A}$ and $\mathcal{D}$ as follows:
$w_{k}\sim\mathcal{W}=\operatorname{Unif}(\{-1,1\}),\ b_{k}\sim\mathcal{B},\ a_{k}\sim\frac{1}{\sqrt{n}}\mathcal{A},\ k=1,\ldots,n;\quad d\sim\mathcal{D}.$ (2)
We assume $\mathcal{W}$ is uniform on the set $\{-1,1\}$, i.e., $P(\mathcal{W}=1)=P(\mathcal{W}=-1)=1/2$, and $\mathcal{B}$, $\mathcal{A}$, $\mathcal{D}$ are sub-Gaussian random variables. Further, assume that $\mathcal{B}$ has a continuous density function $p_{\mathcal{B}}$ and that the random vector $(\mathcal{W},\mathcal{B})$ is symmetric, i.e., $(\mathcal{W},\mathcal{B})$ and $(-\mathcal{W},-\mathcal{B})$ have the same distribution. We use $\mu$ for the distribution of $(\mathcal{W},\mathcal{B})$, and $\hat{\theta}=\operatorname{vec}(\hat{\mathbf{w}},\hat{\mathbf{b}},\hat{\boldsymbol{a}},\hat{d})$ for the initial parameter vector.
@@
Consider a finite training data set $\{(x_{i},y_{i})\}_{i=1}^{m}\subset\mathbb{R}\times\mathbb{R}$. Consider the mean squared error $L(\theta)=\frac{1}{2m}\sum_{i=1}^{m}(f(x_{i},\theta)-y_{i})^{2}$. Given a strictly convex and twice differentiable potential function $\Phi$ on the parameter space, the Bregman divergence (Bregman, 1967) with respect to $\Phi$ is defined as $D_{\Phi}(\theta,\theta^{\prime})=\Phi(\theta)-\Phi(\theta^{\prime})-\langle\nabla\Phi(\theta^{\prime}),\theta-\theta^{\prime}\rangle$. The mirror descent parameter update (Nemirovskij & Yudin, 1983; Beck & Teboulle, 2003) for minimizing the loss $L(\theta)$ is defined by:
@@
The dynamics for the predictions can be given as follows. Let $\mathbf{y}=[y_{1},\ldots,y_{m}]^{T}\in\mathbb{R}^{m}$ denote the vector of the training labels and $\mathbf{f}(t)$ denote the vector of the network prediction on the training input data points at time $t$, i.e., $\mathbf{f}(t)=[f(x_{1},\theta(t)),\ldots,f(x_{m},\theta(t))]^{T}\in\mathbb{R}^{m}$. By the chain rule,
$\frac{\mathrm{d}}{\mathrm{dt}}\mathbf{f}(t)$ $=-n\eta H(t)(\mathbf{f}(t)-\mathbf{y}),$ (5)
where $\text{where }H(t)$ $\triangleq n^{-1}J_{\theta}\mathbf{f}(t)(\nabla^{2}\Phi(\theta(t)))^{-1}J_{\theta}(\mathbf{f}(t))^{T}.$
Here $J_{\theta}\mathbf{f}(t)=[\nabla_{\theta}f(x_{1},\theta(t)),\ldots,\nabla_{\theta}f(x_{m},\theta(t))]^{T}\in\mathbb{R}^{m\times(3n+1)}$ denotes the Jacobian matrix of the network output $\mathbf{f}(t)$ with respect to the network parameter $\theta$. We see in (5) that different choices of $\Phi$ induce different modifications of the Gram matrix $J_{\theta}\mathbf{f}(t)(J_{\theta}\mathbf{f}(t))^{T}$.
@@ 3.1 Implicit bias of mirror flow with unscaled potentials
The potential function $\Phi\colon\mathbb{R}^{3n+1}\rightarrow\mathbb{R}$ satisfies: (i) $\Phi$ can be written in the form $\Phi(\theta)=\sum_{k=1}^{3n+1}\phi(\theta_{k}-\hat{\theta}_{k})$, where $\phi$ is a real-valued function on $\mathbb{R}$; (ii) $\Phi$ is twice continuously differentiable; (iii) $\nabla^{2}\Phi(\theta)$ is positive definite for all $\theta\in\mathbb{R}^{3n+1}$.
@@
The following theorem gives a function space description of the implicit bias of mirror flow for univariate least squares regression with wide neural networks with ReLU activation $\sigma(s)=\max\{0,s\}$.
@@
Consider a two layer ReLU network (1) with $n$ hidden units, where we assume $n$ is sufficiently large. Consider parameter initialization (2) and, specifically, let $p_{\mathcal{B}}(b)$ denote the density function for random variable $\mathcal{B}$. Consider any finite training data set $\{(x_{i},y_{i})\}_{i=1}^{m}$ that satisfies $x_{i}\neq x_{j}$ when $i\neq j$ and $\{x_{i}\}_{i=1}^{M}\subset\operatorname{supp}(p_{\mathcal{B}})$, and consider mirror flow (4) with a potential function $\Phi$ satisfying Assumption 1 and learning rate $\eta=\Theta(1/n)$. Then, with high probability over the random parameter initialization, there exist constants $C_{1},C_{2}>0$ such that for any $t\geq 0$,
$\|\theta(t)-\hat{\theta}\|_{\infty}\leq C_{1}n^{-1/2},\ \lim_{n\rightarrow\infty}\|H(t)-H(0)\|_{2}=0,\ \|\mathbf{f}(t)-\mathbf{y}\|_{2}\leq e^{-\eta_{0}C_{2}t}\|\mathbf{f}(0)-\mathbf{y}\|_{2}.$ (6)
Moreover, letting $\theta(\infty)=\lim_{t\rightarrow\infty}\theta(t)$, we have for any given $x\in\mathbb{R}$, that $\lim_{n\rightarrow\infty}|f(x,\theta(\infty))-f(x,\hat{\theta})-\bar{h}(x)|=0$, where $\bar{h}(\cdot)$ is the solution to the following variational problem:
@@
To interpret the result, (6) means that mirror flow exhibits lazy training and the network behaves like a kernel model with kernel $H(0)$. Interestingly, the choice of the potential $\Phi$ has no effect on the solution to (7), which implies that mirror flow has the same implicit bias as ordinary gradient flow. The reason behind this is that, loosely speaking, since the parameters stay nearly unchanged during training, the effect of the potential reduces to the effect of a quadratic form, i.e., $\Phi(\theta)\simeq c\|\theta\|^{2}_{2}$ for some constant $c>0$. Further below we discuss scaled potentials breaking outside this regime.
To interpret the implicit bias described in (7), we discuss the effect of the different functionals on the solution function: (1) $\mathcal{G}_{1}$ serves as a curvature penalty, which can be viewed as the $p_{\mathcal{B}}^{-1}$-weighted $L_{2}$ norm of the second derivative of $h$. Note in our setting $p_{\mathcal{B}}$ corresponds to the density of $\mathcal{B}/\mathcal{W}$, which is the breakpoint of a ReLU in the network. (2) $\mathcal{G}_{2}$ favors opposite slopes at positive and negative infinity. (3) The effect of $\mathcal{G}_{3}$ is less clear under general sub-Gaussian initializations. In the case where $\mathcal{B}$ is compactly supported, $\mathcal{G}_{3}$ can be interpreted as follows: if $h$ is generally above the $x$-axis (specifically, $h(B)+h(-B)>0$), then $\mathcal{G}_{3}$ encourages $h^{\prime}(+\infty)>h^{\prime}(-\infty)$ and thus favors a U-shaped function; if on the contrary $h$ is below the $x$-axis, $\mathcal{G}_{3}$ favors an inverted U-shaped function.
The functional $\mathcal{G}_{1}$ is familiar from the literature on space adaptive interpolating splines and in particular natural cubic splines when $p_{\mathcal{B}}$ is constant (see, e.g., Ahlberg et al., 2016). However, with the addition of $\mathcal{G}_{2}$ and $\mathcal{G}_{3}$, the solution to (7) has regularized boundary conditions. In Figure 1, we illustrate numerically that gradient flow returns natural cubic splines only on very specific training data. This aligns with the observation made by Jin & Montúfar (2023) that gradient descent returns natural cubic splines when some unknown linear adjustment is applied to the training data.
@@
As observed by Li et al. (2022), mirror flow is closely related to gradient flow with commuting reparameterization. In Appendix D, we use our results to characterize the implicit bias of gradient descent with particular reparametrizations.
@@
For networks with absolute value activation $\sigma(s)=|s|$, a similar result to Theorem 1 holds, which we formally present in Theorem 23. The only change is that $\mathcal{G}_{2}$ and $\mathcal{G}_{2}$ disappear in the objective functional. In fact, in Proposition 32 we show that absolute value networks form a proper subspace of ReLU networks where both $\mathcal{G}_{2}$ and $\mathcal{G}_{3}$ vanish.
@@
By Theorem 2, in the infinite width limit, mirror flow with hypentropy potential $\rho_{\beta}$ for any $\beta>0$ has the same bias as ordinary gradient flow. Taking $\beta\rightarrow 0^{+}$, we recover a result by Zhang et al. (2019) stating NGD converges to same point as gradient descent.

### v2

# Implicit Bias of Mirror Flow for Shallow Neural Networks in Univariate Regression

@@ Abstract
We examine the implicit bias of mirror flow in least squares error regression with wide and shallow neural networks. For a broad class of potential functions, we show that mirror flow exhibits lazy training and has the same implicit bias as ordinary gradient flow when the network width tends to infinity. For univariate ReLU networks, we characterize this bias through a variational problem in function space. Our analysis includes prior results for ordinary gradient flow as a special case and lifts limitations which required either an intractable adjustment of the training data or networks with skip connections. We further introduce scaled potentials and show that for these, mirror flow still exhibits lazy training but is not in the kernel regime. For univariate networks with absolute value activations, we show that mirror flow with scaled potentials induces a rich class of biases, which generally cannot be captured by an RKHS norm. A takeaway is that whereas the parameter initialization determines how strongly the curvature of the learned function is penalized at different locations of the input space, the scaled potential determines how the different magnitudes of the curvature are penalized.
@@ 1 Introduction
Mirror descent is a broad class of first-order optimization algorithms that generalize ordinary gradient descent (Nemirovskij & Yudin, 1983). It can be viewed as using a general distance-like function, defined by the choice of a convex potential function, instead of the usual Euclidean squared distance to determine the update direction in the search space (Beck & Teboulle, 2003). The choice of the geometry of the parameter space of a learning system is crucial, as it can affect the speed of convergence and the implicit bias of parameter optimization (Neyshabur et al., 2017). The implicit bias of mirror descent has been studied theoretically for linear models (Gunasekar et al., 2018a; Sun et al., 2023; Pesme et al., 2024), matrix sensing (Wu & Rebeschini, 2021), and also for nonlinear models at least in a local sense (Azizan et al., 2022). However, we are not aware of any works characterizing the implicit bias of mirror decent for neural networks describing more than the possible properties of the parameters also the nature of the returned solution functions. This is the gap that we target in this article. We focus on the particular setting of mirror flow applied to least squares error regression with shallow and infinitely wide neural networks.
@@
We show for networks with standard parametrization that, under a broad class of potential functions, mirror flow displays lazy training, meaning that even though the loss converges to zero the parameters remain nearly unchanged. We show in turn that mirror flow has the same implicit bias as gradient flow when the network width tends to infinity. This is surprising because, in contrast, for finite-dimensional parameter models mirror flow has been observed to exhibit a different implicit bias than gradient descent (see, e.g., Gunasekar et al., 2018a; Sun et al., 2023). We then proceed to characterize the particular solution function that is returned by mirror flow for wide ReLU networks, by expressing the difference between the network functions before and after training, $g(x)=f_{\mathrm{trained}}(x)-f_{\mathrm{initial}}(x)$, as the solution to a variational problem. By examining the function representation cost, we show that mirror flow is biased towards minimizing a weighted $L^{2}$ norm of the second-order derivative of the difference function, $\int p^{-1}(x)(g^{\prime\prime}(x))^{2}\mathrm{d}x$, where the weight $p^{-1}$ is determined by the network parameter initialization, and two additional terms that control the first-order derivative. This includes as a special case a previous result for gradient flow and ReLU activations by Jin & Montúfar (2023). Our analysis also removes a key limitation of that result, which required either an intractable linear adjustment of the training data or otherwise adding skip connections to the network. We observe that other previous works on the implicit bias of gradient descent suggested that shallow ReLU networks in the kernel regime could be well-approximated by a cubic spline interpolator of the training data (Williams et al., 2019; Heiss et al., 2019; Sahs et al., 2022) but overlooked the need of linearly adjusting the training data or adding skip connections. We illustrate that even a slight modification of the training data by applying a constant shift to all training labels can lead to significantly different solutions upon training which are biased not only in terms of curvature but also slopes.
@@ 1.1 Contributions
The goal of this article is to provide insights into the implicit bias of mirror flow for shallow and wide neural networks with standard parametrization. Our contributions can be summarized as follows.
  • We show for a broad class of potential functions that, mirror flow, when applied to networks with general input dimension, displays lazy training. We show that in this case the implicit bias of mirror flow is equivalent to that of gradient flow when the number of hidden units tends to infinity.
  • We characterize the implicit bias of mirror flow for univariate networks in function space. Concretely, we express the difference of the trained and initialized network function as the solution to a variational minimization problem in function space (Theorem 2).
  • We introduce a class of scaled potentials for which we show that mirror flow, applied to networks with general input dimension, falls in the lazy regime but not in the kernel regime. We characterize the corresponding implicit bias for infinitely wide networks and observe it is strongly dependent on the potential and thus different from ordinary gradient flow.
  • Finally, we characterize the implicit bias of mirror flow with scaled potentials in function space for univariate networks with absolute value activation function (Theorem 9). Our results show that the choice of the potential biases the magnitude of the curvature of the solution function.
@@ 1.2 Related works
The implicit bias of mirror descent has been studied for under-determined linear models. In classification problems with linearly separable data, Sun et al. (2022) show that mirror descent with homogeneous potentials converges in direction to a max-margin classifier with respect to a norm defined by the potential. Pesme et al. (2024) extend this work to non-homogeneous potentials. For regression, Gunasekar et al. (2018a); Azizan & Hassibi (2019) show that mirror descent is biased towards the minimizer of the loss that is closest to the initialization with respect to the Bregman divergence defined by the potential. Azizan et al. (2022) extend this result to nonlinear models, albeit placing a strong assumption that the initial parameter is sufficiently close to a solution. For matrix sensing, Wu & Rebeschini (2021) show mirror descent with spectral hypentropy potential tends to recover low-rank matrices. Li et al. (2022) connect gradient flow with commuting reparameterization to mirror flow, and show that both have identical trajectories and implicit biases. This connection is further studied by Chou et al. (2023) specifically for linear models.
@@
For a parameterized model and a given cost function, e.g., a norm, on the parameter space, the representation cost of a target function refers to the minimal cost required to implement that function. Existing work has revealed a close connection between representation cost and the implicit bias of gradient-based optimization methods, as the latter can often be described by specific parameter norm minimization problems. A number of works have studied the representation cost for neural networks with varying architectures (Gunasekar et al., 2018b; Savarese et al., 2019; Ongie et al., 2020; Boursier & Flammarion, 2023; Jacot, 2023; Parkinson et al., 2024). Among these, Savarese et al. (2019); Ongie et al. (2020); Boursier & Flammarion (2023) have focused on infinitely wide and shallow ReLU networks, similar to our work. While these works focused on the $L^{1}$ norm on the output weights as the cost function, we considered the $L^{2}$ norm and encompasses a broad family of cost functions induced by Bregman divergence. Moreover, instead of studying representation cost in a static and optimization-free manner, we provide detailed analyses of the optimization dynamics of mirror flow and characterize the implicit bias with concrete convergence guarantees.
@@ 2 Problem Setup
We focus on fully-connected two-layer neural networks with $d$ input units, $n$ hidden units and one output unit, defined by:
$f(x,\theta)=\sum_{k=1}^{n}a_{k}\sigma(w_{k}^{T}x-b_{k})+d,$ (1)
where $\sigma(z)$ is a nonlinear activation function. We write $\theta=\operatorname{vec}(\boldsymbol{W},\boldsymbol{b},\boldsymbol{a},d)\in\mathbb{R}^{p}$, with $p=n(d+2)+1$, for the vector of all parameters, comprising the input weights $\boldsymbol{W}=(w_{1},\ldots,w_{n})^{T}\in\mathbb{R}^{n\times d}$, input biases $\boldsymbol{b}=(b_{1},\ldots,b_{n})^{T}\in\mathbb{R}^{n}$, output weights $\boldsymbol{a}=(a_{1},\ldots,a_{n})^{T}\in\mathbb{R}^{n}$, and output bias $d\in\mathbb{R}$. We assume the parameters of the network are initialized by independent samples of mutually independent random variables $\mathcal{W},\mathcal{B},\mathcal{A}$ and $\mathcal{D}$ as follows:
$w_{k}\sim\mathcal{W}=\operatorname{Unif}(\mathbb{S}^{d-1}),\ b_{k}\sim\mathcal{B},\ a_{k}\sim\frac{1}{\sqrt{n}}\mathcal{A},\ k=1,\ldots,n;\quad d\sim\mathcal{D}.$ (2)
We assume $\mathcal{W}$ is uniform on the unit sphere in $\mathbb{R}^{d}$, and $\mathcal{B}$, $\mathcal{A}$, $\mathcal{D}$ are sub-Gaussian random variables. Further, assume that $\mathcal{B}$ has a continuous density function $p_{\mathcal{B}}$ and that the random vector $(\mathcal{W},\mathcal{B})$ is symmetric, i.e., $(\mathcal{W},\mathcal{B})$ and $(-\mathcal{W},-\mathcal{B})$ have the same distribution. We use $\mu$ for the distribution of $(\mathcal{W},\mathcal{B})$, and $\hat{\theta}=\operatorname{vec}(\hat{\boldsymbol{W}},\hat{\boldsymbol{b}},\hat{\boldsymbol{a}},\hat{d})$ for the initial parameter vector.
@@
Consider a finite training data set $\{(x_{i},y_{i})\}_{i=1}^{m}\subset\mathbb{R}^{d}\times\mathbb{R}$. Consider the mean squared error $L(\theta)=\frac{1}{2m}\sum_{i=1}^{m}(f(x_{i},\theta)-y_{i})^{2}$. Given a strictly convex and twice differentiable potential function $\Phi$ on the parameter space, the Bregman divergence (Bregman, 1967) with respect to $\Phi$ is defined as $D_{\Phi}(\theta,\theta^{\prime})=\Phi(\theta)-\Phi(\theta^{\prime})-\langle\nabla\Phi(\theta^{\prime}),\theta-\theta^{\prime}\rangle$. The mirror descent parameter update (Nemirovskij & Yudin, 1983; Beck & Teboulle, 2003) for minimizing the loss $L(\theta)$ is defined by:
@@
The dynamics for the predictions can be given as follows. Let $\boldsymbol{y}=[y_{1},\ldots,y_{m}]^{T}\in\mathbb{R}^{m}$ denote the vector of the training labels and $\boldsymbol{f}(t)$ denote the vector of the network prediction on the training input data points at time $t$, i.e., $\boldsymbol{f}(t)=[f(x_{1},\theta(t)),\ldots,f(x_{m},\theta(t))]^{T}\in\mathbb{R}^{m}$. By the chain rule,
$\frac{\mathrm{d}}{\mathrm{dt}}\boldsymbol{f}(t)$ $=-n\eta H(t)(\boldsymbol{f}(t)-\boldsymbol{y}),$ (5)
where $\text{where }H(t)$ $\triangleq n^{-1}J_{\theta}\boldsymbol{f}(t)(\nabla^{2}\Phi(\theta(t)))^{-1}J_{\theta}(\boldsymbol{f}(t))^{T}.$
Here $J_{\theta}\boldsymbol{f}(t)=[\nabla_{\theta}f(x_{1},\theta(t)),\ldots,\nabla_{\theta}f(x_{m},\theta(t))]^{T}\in\mathbb{R}^{m\times p}$ denotes the Jacobian matrix of the network output $\boldsymbol{f}(t)$ with respect to the network parameter $\theta$. We see in (5) that different choices of $\Phi$ induce different modifications of the Gram matrix $J_{\theta}\boldsymbol{f}(t)(J_{\theta}\boldsymbol{f}(t))^{T}$.
@@ 3.1 Implicit bias of mirror flow with unscaled potentials
The potential function $\Phi\colon\mathbb{R}^{p}\rightarrow\mathbb{R}$ satisfies: (i) $\Phi$ can be written in the form $\Phi(\theta)=\sum_{k=1}^{p}\phi(\theta_{k}-\hat{\theta}_{k})$, where $\phi$ is a real-valued function on $\mathbb{R}$; (ii) $\Phi$ is twice continuously differentiable; (iii) $\nabla^{2}\Phi(\theta)$ is positive definite for all $\theta\in\mathbb{R}^{p}$.
@@
The following theorem characterizes the implicit bias of mirror flow for least squares regression with wide neural networks with ReLU activation $\sigma(s)=\max\{0,s\}$.
@@
Consider a two layer ReLU network (1) with $d\geq 1$ input units and $n$ hidden units, where we assume $n$ is sufficiently large. Consider parameter initialization (2) and, specifically, let $p_{\mathcal{B}}(b)$ denote the density function for random variable $\mathcal{B}$. Consider any finite training data set $\{(x_{i},y_{i})\}_{i=1}^{m}$ that satisfies $x_{i}\neq x_{j}$ when $i\neq j$ and $\{\|x_{i}\|_{2}\}_{i=1}^{m}\subset\operatorname{supp}(p_{\mathcal{B}})$, and consider mirror flow (4) with a potential function $\Phi$ satisfying Assumption 1 and learning rate $\eta=\Theta(1/n)$. Then, there exist constants $C_{1},C_{2}>0$ such that with high probability over the random parameter initialization, for any $t\geq 0$,
$\|\theta(t)-\hat{\theta}\|_{\infty}\leq C_{1}n^{-1/2},\ \lim_{n\rightarrow\infty}\|H(t)-H(0)\|_{2}=0,\ \|\boldsymbol{f}(t)-\boldsymbol{y}\|_{2}\leq e^{-\eta_{0}C_{2}t}\|\boldsymbol{f}(0)-\boldsymbol{y}\|_{2}.$ (6)
Moreover, letting $\theta(\infty)=\lim_{t\rightarrow\infty}\theta(t)$, we have for any given $x\in\mathbb{R}^{d}$, that $\lim_{n\rightarrow\infty}|f(x,\theta(\infty))-f(x,\theta_{\mathrm{GF}}(\infty))|=0$, where $\theta_{\mathrm{GF}}(\infty)$ denotes the limiting point of gradient flow, i.e., mirror flow (4) with $\Phi(\theta)=\|\theta\|_{2}^{2}$, on the same training data and initial parameter.
Assuming univariate input data, i.e., $d=1$, we have for any given $x\in\mathbb{R}$, that $\lim_{n\rightarrow\infty}|f(x,\theta(\infty))-f(x,\hat{\theta})-\bar{h}(x)|=0$, where $\bar{h}(\cdot)$ is the solution to the following variational problem:
@@
To interpret the result, (6) means that mirror flow exhibits lazy training and the network behaves like a kernel model with kernel $H(0)$. The theorem indicates that mirror flow has the same implicit bias as ordinary gradient flow, and the reason behind this is that, loosely speaking, since the parameters stay nearly unchanged during training, the effect of the potential reduces to the effect of a quadratic form, i.e., $\Phi(\theta)\simeq\|\theta\|^{2}_{2}$. Further below we discuss scaled potentials breaking outside this regime.
To interpret the implicit bias described in (7), we discuss the effect of the different functionals on the solution function: (1) $\mathcal{G}_{1}$ serves as a curvature penalty, which can be viewed as the $p_{\mathcal{B}}^{-1}$-weighted $L^{2}$ norm of the second derivative of $h$. Note in our setting $p_{\mathcal{B}}$ corresponds to the density of $\mathcal{B}/\mathcal{W}$, which is the breakpoint of a ReLU in the network. (2) $\mathcal{G}_{2}$ favors opposite slopes at positive and negative infinity. (3) The effect of $\mathcal{G}_{3}$ is less clear under general sub-Gaussian initializations. In the case where $\mathcal{B}$ is compactly supported, $\mathcal{G}_{3}$ can be interpreted as follows: if $h$ is generally above the $x$-axis (specifically, $h(B)+h(-B)>0$), then $\mathcal{G}_{3}$ encourages $h^{\prime}(+\infty)>h^{\prime}(-\infty)$ and thus favors a U-shaped function; if on the contrary $h$ is below the $x$-axis, $\mathcal{G}_{3}$ favors an inverted U-shaped function.
The functional $\mathcal{G}_{1}$ is familiar from the literature on space adaptive interpolating splines and in particular natural cubic splines when $p_{\mathcal{B}}$ is constant (see, e.g., Ahlberg et al., 2016). However, with $\mathcal{G}_{2}$ and $\mathcal{G}_{3}$, the solution to (7) has regularized boundary conditions. In Figure 1, we illustrate numerically that gradient flow returns natural cubic splines only on very specific training data.
Remark 3 (Relaxed potential assumption).
Assumption 1 requires the potential $\Phi$ to be “centered” at the initial parameter $\hat{\theta}$. In Appendix A.4, we show Theorem 2 also holds for potentials of the form $\Phi(\theta)=\sum_{k}\phi(\theta_{k})$, given the initial input biases are sampled from a bounded distribution.
@@
As observed by Li et al. (2022), mirror flow is equivalent to gradient flow with commuting reparameterization. In Appendix D, we use our results to characterize the implicit bias of gradient flow with particular reparametrizations.
@@
For networks with absolute value activation $\sigma(s)=|s|$, a similar result to Theorem 2 holds, which we formally present in Theorem 24. The only change is that $\mathcal{G}_{2}$ and $\mathcal{G}_{2}$ disappear in the objective functional.
@@
Theorem 2 holds for hypentropy potential $\rho_{\beta}$ for any $\beta>0$, and in particular, in the limit as $\beta\rightarrow 0^{+}$. This is consistent with a result by Zhang et al. (2019) stating NGD converges to the same point as ordinary gradient descent.